OpenStudy (anonymous):

x/x+3 + 2/5 =0 solve and express it in simplest fraction

OpenStudy (anonymous):

Take the LCM of 5 and (x+3).. Can you tell me what is the LCM here..

OpenStudy (anonymous):

Or simply multiply by 5(x+3) to both the sides..

OpenStudy (anonymous):

5x+15

OpenStudy (anonymous):

Need not to multiply them you can multiply them at the end.. So LCM is 5(x+3)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Simply multiply both the sides by 5(x+3) both the sides can you do it ??

OpenStudy (anonymous):

im sorry, im lost....

OpenStudy (anonymous):

\[5(x+3) \times \frac{x}{x+3} + 5(x+3) \times \frac{2}{5} = 0 \times 5(x+3)\] Can you do this ??

OpenStudy (anonymous):

In very first term what is cancelling ?

OpenStudy (anonymous):

x+3

OpenStudy (anonymous):

yes so cancel it.. And what are you left with now??

OpenStudy (anonymous):

7x+6=0???

OpenStudy (anonymous):

oohhhh so x=6/7?

OpenStudy (anonymous):

No minus mistake is there..

OpenStudy (anonymous):

oh its -6/7

OpenStudy (anonymous):

yes well done..

OpenStudy (anonymous):

thaqnk you